OpenStudy (anonymous):

What is the difference between angular speed, angular velocity, and tangential velocity? I'm confused on what each is used for.

OpenStudy (anonymous):

If an object is going a circle, then a radius drawn from the center of the circle ill change its angle (with the horizontal or x direction) at a rate which is the angular velocity vector. The tangential velocity is the change in distance divided by a short time interval, v = ds/dt and it is at right angles to the radius there.

OpenStudy (anonymous):

What about angular speed?

OpenStudy (anonymous):

I think I understand angular velocity. It's just the time it takes for the angle to change?

OpenStudy (anonymous):

yeah angular velocity, omega, is \[ \omega = \frac{\Delta \theta}{\Delta t}\] change in the angle over change in time.

OpenStudy (anonymous):

I'm not familiar with angular speed It sounds like angular velocity as a scalar rather than a vector.

OpenStudy (anonymous):

Scalar?

OpenStudy (anonymous):

a scalar is something that has a magnitude but no direction. ie, if you're running at 10 m/s, and you're running North Your speed , a scalar, is just the value 10 m/s \[u=10m/s\] whereas the velocity, a vector, is 10 m/s North \[\vec{v}=10m/s \ \ \text{North}\]